OpenStudy (anonymous):

The sum of three consecutive even integers is –114. What is the least of these numbers?

OpenStudy (anonymous):

Let the consecutive even integers be x-2.x,x+2 x-2+x+x+2=-114 3x=-114 x=-114/3=-38 find the least